The sum of the ages of Tony and Teddy equals thrice their mother’s age. Tony is two years younger than Teddy. If the age of the mother is 42, find the ages of Tony and Teddy.
I've done:
t + d = 3m
t = d - 2
m = 42
I "plugged" in the numbers but the numbers are really high at the end. What am I doing wrong here?
Thank you.

Answer by stanbon(75887)

The sum of the ages of Tony and Teddy equals thrice their mother’s age. Tony is two years younger than Teddy. If the age of the mother is 42, find the ages of Tony and Teddy.
I've done:
t + d = 3m
t = d - 2
m = 42
--------
So:
t + d = 3*42 = 126
t - d = -2
-----
Add to get:
2t = 124
Tony = 62 years
----
Solve for "d":
d = t + 2
Teddy = 64 years
====================
Cheers,
Stan H.
Note: The answers are unrealistic
but mathematically correct.
Not a good problem.